an investment pays 7% compound semiannually. You want to have $ 19000.00 in the future , how much should you deposit now to have that amount in 5 years from now.
Answer by rfer(16322)

19000=P(1+0.07/2)^5*2
19000=P(1.035)^10
19000=P(1.41)
19000/1.41=P
$13.475=P